# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ Acknowledgments
We thank Karin Jacobs and Christian Spengler for providing their data of nanorough surfaces, and Jens Uwe Neurohr for Figure 5. This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265, under grant numbers HU 1874/5-1, ME 1361/16-1, KL 3391/2-2, WI 5527/1-1, and LO 418/25-1, as well as by the Helmholtz Association and the DLR via the Helmholtz Young Investigator Group “DataMat”.
@@ Appendix A Further simulation results
  • Table 11: Simulation results for the 2-dimensional rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$
@@
  • Table 13: Simulation results for the 3-dimensional rectangle $[-\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$.
Table 11: Results for the 2-dimensional rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=50$ and $R_{n}=2$. We took the average of 50 renditions. All values are rounded to the fourth significant digit.
@@
Table 12: Results for a 2-dimensional rectangle with rounded vertices intersected with a grid of resolution $a=0.005$. Here $R_{r_{0}}$ denotes the parallel set of the rectangle $[-\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ with parameter $r_{0}>0$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 10 renditions.
@@
Table 13: Results for the 3-dimensional rectangle $[-\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 28 renditions.

### v2

@@ Acknowledgments
We thank Karin Jacobs and Christian Spengler for providing their data of nanorough surfaces, and Jens Uwe Neurohr for Figure 12. This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265, under grant numbers HU 1874/5-1, ME 1361/16-1, KL 3391/2-2, WI 5527/1-1, and LO 418/25-1, as well as by the Helmholtz Association and the DLR via the Helmholtz Young Investigator Group “DataMat”. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). The hardware is funded by the German Research Foundation (DFG).
@@ Appendix A Further simulation results
  • Table 11: Simulation results for the 2-dimensional rectangle $\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$
@@
  • Table 13: Simulation results for the 3-dimensional rectangle $\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$.
Table 11: Results for the 2-dimensional rectangle $\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=50$ and $R_{n}=2$. We took the average of 50 renditions. All values are rounded to the fourth significant digit.
@@
Table 12: Results for a 2-dimensional rectangle with rounded vertices intersected with a grid of resolution $a=0.005$. Here $R_{r_{0}}$ denotes the parallel set of the rectangle $\frac{3}{2},\frac{3}{2}]\times[-\frac{5}{2},\frac{5}{2}]$ with parameter $r_{0}>0$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 10 renditions.
@@
Table 13: Results for the 3-dimensional rectangle $\frac{1}{2},\frac{1}{2}]\times[-1,1]\times[-\frac{3}{2},\frac{3}{2}]$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=5$ and $R_{n}=1$. We took the average of 28 renditions.